OpenStudy (anonymous):

What is 35% of 45? A. 10.5 B. 21 C. 22.75 D. 15.75

OpenStudy (anonymous):

will medal or fan

OpenStudy (anonymous):

35% in fraction is: 35/100 = 0.35 Then do 0.35 * 45 = ?

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

15.75?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

If that's what the calculator says, then yea.

OpenStudy (anonymous):

?

OpenStudy (vhtran):

thats correct.

OpenStudy (anonymous):

i just chose the one that was the closest to one of the answers

jhonyy9 (jhonyy9):

45 --- 100% x ---- 35% ------------- x = 35*45/100 = ?

jhonyy9 (jhonyy9):

D is right sure

OpenStudy (vhtran):

0.35 times 45 = 15.75